Sum of complement and supplement angle is 200. What is the measure of the angle
Answer by nerdybill(7384)

Sum of complement and supplement angle is 200. What is the measure of the angle
.
Let x = the angle
then
90-x = complement
180-x = supplement
.
90-x + 180-x = 200
90+180-2x = 200
270-2x = 200
-2x = -70
x = 35 degree